OpenStudy (anonymous):

Solve using the substitution method x+3y=10 xy=7

hero (hero):

What's the matter? You didn't like my method?

OpenStudy (anonymous):

y = 1 and x = 7

OpenStudy (anonymous):

Hmm that's weird lol. Yeah the two solutions that Juanita gave worked now I just have to go through it and make sure I understand it for the exam! Thanks for helping out guys I appreciate it!

hero (hero):

Yeah, I made a calculation mistake. I hate Luis Rivera because he uses programs

hero (hero):

Luis Rivera provided the two solutions, not Juanita by the way. Nevertheless, I found my calculation mistake and my method is still intact.

OpenStudy (anonymous):

Yeah I just caught that lol. I need to prepare because I have to take up to calc 3 for my major and Im not very good at math ha. I just have to keep pushing I guess!

OpenStudy (anonymous):

do you need me to show work

hero (hero):

x + 3y = 10 x = 10 - 3y x = 7/y x = x 10 - 3y = 7/y y(10-3y) = 7 10y - 3y^2 = 7 3y^2 - 10y + 7 = 0 3y^2-3y-7y+7 = 0 3y(y-1)-7(y-1)=0 (y-1)(3y-7)=0 y = 1 y = 3/7 x = 7 x = 3

hero (hero):

There's my corrected solution

hero (hero):

Juanita, there truly are two solutions

OpenStudy (anonymous):

hero -- your 3rd line is incorrect - the solutions i gave are correct - you can not add 10-3y and get 7/7

hero (hero):

juanita, do I need to put this in writing for you? nobody added 10-3y to get 7/7

hero (hero):

You should plug the solutions back in and see what you get. For some reason, you're convinced your answer is correct when you have not even posted a solution

OpenStudy (anonymous):

Hero x + 3y = 10 x = 10 - 3y x = 7/y your 3rd line is incorrect - i meant to type 7/Y

hero (hero):

how is that incorrect? Explain

hero (hero):

Don't give me the bit about dividing by a variable because I multiply it back in

hero (hero):

Which is totally valid

OpenStudy (anonymous):

how can you combine 10-3y to get 7/y? that is not proper nor correct

hero (hero):

You don't understand well enough, it is okay though.

hero (hero):

It's called substitution by the way

hero (hero):

Two expressions equal x: x = 10 - 3y x = 7/y Now we set x = x meaning those two expressions are set equal: x = x 10 - 3y = 7/y If you don't understand that, I can't help you.

OpenStudy (anonymous):

x+3y = 10 xy = 7 x = 10-3y (10-3y)(y) = 7 10y -3y^2 = 7 using quadratic equation zeros are x = 7/3 and 1 subbing back in to xy = 7 y = 3/7 and 7 those are the actually correct answers.

OpenStudy (anonymous):

sweetheart i was just telling you that your third line was incorrect. Have a great day

hero (hero):

Actually, you didn't. Something is wrong with your answers.

hero (hero):

The points are (7,1) and (3,3/7)

hero (hero):

Your approach is incorrect. You either inputted it wrong or something.

hero (hero):

Juanita, if you can find someone who agrees the my solution is not correct, I'll give you the respect you're looking for.

hero (hero):

I see why you're confused now Juanita. I'll show you why my third line is correct: \[xy = 7\] \[\frac{xy}{y} = \frac{7}{y}\] \[x = \frac{7}{y}\] x = 7/y is my third line. These are the steps to get to the third line.